The width of a rectangle is 5 inches less than its length. If the area of the rectangle is 266 square inches, what is its widthH?
10 in
19 in
37 in
14 in

Answer by lwsshak3(11628)

The width of a rectangle is 5 inches less than its length. If the area of the rectangle is 266 square inches, what is its widthH?
let x=length of rectangle
x-5=width of rectangle
length*width=area=266
x(x-5)=266
x^2-5x-266=0
(x-14)(x+19)=0
x=14
x-5=9
=width of rectangle=9 inches